OpenStudy (anonymous):

Determine whether the graphs of the following equations are parallel or perpendicular. Explain. 3x-9y=9, 3y=x+12 ,2x-6y=12

OpenStudy (jdoe0001):

solve each of them for "y", what do you get?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=8j3kRdiRjPA

OpenStudy (anonymous):

just bring everything that is not y to the right.. so u get -9y=-3x+9... then do simple algebra so that its not -9y but +y

OpenStudy (anonymous):

Im so confused on the whole problem. i don't understand this at all

OpenStudy (anonymous):

well we are trying to find the slope of the equation... and one of the simplest ways to do that when you have a equation of a line is to look at the number before the 'x' when the equation is in terms of 'y'.

OpenStudy (anonymous):

-9y= -3x +9 9y = 3x + 9 y = \[\frac{ 3x }{ 9 } + \frac{ 9 }{ 9 }\]

OpenStudy (anonymous):

get it so far?

OpenStudy (anonymous):

not really the starting equation is 3x-9y=9

OpenStudy (anonymous):

It'll help you tons if you get Y alone. on all 3 of your equations. Then check if they have the same slope. All parallel lines have the same slope.

OpenStudy (anonymous):

How do you get y by itself on 3y=x+12

OpenStudy (anonymous):

you divide 3 on both sides. so, divide 3/x and 12/3 you'll get y=1/3+4.

OpenStudy (anonymous):

wouldn't it be parallel because when i made all the equations y= i got 1/3 as the slope on them all is that right

OpenStudy (anonymous):

if for all equations your slope was 1/3 then yes, it's correct and they're all parallel.

OpenStudy (anonymous):

okay thanks